OpenStudy (anonymous):

Let n be a natural number. Which of the following equals i^4n? a) -1 b) i c) 1 d) -1 If somebody could walk me through this, I will award a medal!

OpenStudy (phi):

do you mean \[ i^{4n} \] ? if so , write it as \[ \left(i^4\right)^n \]

OpenStudy (phi):

that uses the rule that \[ \left(a^b\right)^c = a^{bc}\] in reverse.

OpenStudy (phi):

now figure out i^4 i*i * i*i i*i is what ?

OpenStudy (anonymous):

Okay, I think I understand now but I'd really appreciate someone clarifying for me. Since the exponents of i circle around between i, -i, 1, -1 it just depends on the multiple? So i^4n is the same as i^4, which is equal to 1. Yes sorry for the confusion, I mean \[(i^4)^n\]

OpenStudy (phi):

i^4 is 1 so your problem is 1^n where n is a natural number (an integer 1 or bigger) 1^n is 1*1*1*...*1 for n ones obviously that is just 1

OpenStudy (anonymous):

Here, I don't think I wrote it out quite right - it's really meaning i^(4n) Which I can't seem to make open study write correctly. Maybe that means the same thing, because the answer to both is 1.

OpenStudy (phi):

notice we can't play the same trick for i^(3n) writing it as (i^3)^n i*i*i is -i and (-i)^n is (-1)^n * i^n and unless we know more about n, we can't do anything.

OpenStudy (phi):

Do you know about this rule of exponents: \[ \left(a^b\right)^c = a^{bc} \] that is what lets you rewrite i^(4n) as (i^4)^n

OpenStudy (anonymous):

Thank you, that explanation was really helpful. It is 1, but not for the reason I thought (: thanks for the refresher!